OpenStudy (anonymous):

Solve for n: -8 = 4n - 12

OpenStudy (mkmkasim):

n=1

OpenStudy (anonymous):

add 8 to both sides -8+8=4n-12+8

OpenStudy (mkmkasim):

move the 12 over by adding it, that gets you 4=4n. then, divide by 4 to give you n=1

OpenStudy (anonymous):

4n-12+8=0 4n-4=0 4n=4 n=1